OpenStudy (spareb665):

Express ln√ab/c^3 as a sums and difference of logarithms.

OpenStudy (spareb665):



OpenStudy (zehanz):

First, remember: \[\sqrt{x}=x^{\frac{1}{2}}\] Then you'll need a few logarithm rules for this: 1.\[\ln x^k=k \cdot \ln x\] 2.\[\ln ab = \ln a + \ln b\] 3.\[\ln \frac{a}{b}= \ln a - \ln b\]

OpenStudy (zehanz):

If you carefully apply these rules, you'll get the right answer!

OpenStudy (spareb665):

-_-

Parth (parthkohli):

Why the face? I see nothing wrong with the solution.

OpenStudy (zehanz):

Maybe you need a little help with the first step? Here it is:\[\ln \sqrt{\frac{ ab }{ c^3 }}=\ln \left( \frac{ ab }{ c^3 } \right)^{\frac{ 1 }{ 2 }}=\frac{ 1 }{ 2 }\ln \frac{ ab }{ c^3 }\] See? I used my first remark about the square root. Then I applied Rule #1. Now you do the next step! It isn't that hard...

OpenStudy (spareb665):

= 1/2 in ab/c^3 ? using the power rule?

OpenStudy (zehanz):

Yes, I used the power rule to get that. Now go on with the other rules...

OpenStudy (spareb665):

1/2(in ab-in c^3) using the quotient rule?

OpenStudy (zehanz):

Yes! BTW replace "in" with "ln" because it's the natural logarithm (ln)

OpenStudy (spareb665):

ok

OpenStudy (zehanz):

But there is more...

OpenStudy (spareb665):

= 1/2 (In a + In b - 3 In c) using the product and power rules?

OpenStudy (zehanz):

Right. Well done!

OpenStudy (spareb665):

thats it?

OpenStudy (zehanz):

Yup, nothing more to be done!